The cmtt^{∗} package

Mark Wooding

25 May 1996

Contents

1 Introductory note 1

2 Implementation 3 2.1 The package . . . . . . . . 3 2.2 The font definition file . . 5 2.3 The encoding definitions

file . . . . . . . . . . . . . 6

A The GNU General Public Licence 7 A.1 Preamble . . . . . . . . . 7 A.2 Terms and conditions for copying, distribution and modification . . . . . . . . 8

A.3 Appendix: How to Apply These Terms to Your New Programs . . . . . . 12

1 Introductory note

L^{A}TEX has a rather cunning encoding handling system, which makes funny commands like accents work properly independent of the current font’s actual layout. While this works rather well most of the time, the standard tt font has been rather left out of things. L^{A}TEX assumes that the Computer Modern Typewriter fonts have exactly the same layout as the more normal Computer Modern Roman family (i.e., that both conform to the OT1 encoding). This plainly isn’t true, since the Typewriter font contains a bunch of standard ASCII characters which are omitted from the standard Computer Modern fonts, such as curly braces { and }, and the backslash \; these are usually dug up from the maths fonts, which looks fine in normal text, but looks really odd in monospace text. Compare ‘\begin{document}’

to ‘\begin{document}’, for example.

There are two possibilities for dealing with this problem. One is to use the \verb command, which works since all the extra characters in the Typewriter font are in the correct places, or use the DC fonts, which have a proper encoding set up which contains all of these special characters anyway.

Neither of these solutions is perfect. Using \verb causes all manner of little niggly problems: you can’t use it in footnotes or section headings, for example. (There are of course workarounds for this sort of thing: the author’s footnote package provides a footnote environment which will allow verbatim text, and verbatim text in section headings can be achieved if one is sufficiently TEXnical.)

∗The cmtt package is currently at version 1.1, dated 25 May 1996.

Using the DC fonts is fine, although you actually lose a glyph or two. As far as the author is aware, the character ' (an ‘unsexed’ single quote) is not present in the T1-encoded version of Computer Modern Typewriter, although it is hidden away in the original version. The author has found a need for this character in computer listings, and was horrified to discover that it was replaced by a German single quote character (\quotesinglbase).

This package defines a special encoding for the Computer Modern Typewriter font, so that documents can take advantage of its ASCII characters without resorting to verbatim text. (The main advantage of the DC fonts, that words containing accents can be hyohenated, doesn’t really apply to the Typewriter font, since it doesn’t allow hyphenation by default anyway.)

There are several files you’ll need to create:

cmtt.sty tells L^{A}TEX that there’s a new encoding. It also provides some options for customising some aspects of the encoding, and defines some useful commands.

mTTenc.def describes the encoding to L^{A}TEX: it sets up all the appropriate text commands so that they produce beautiful results.

mTTcmtt.fd describes the re-encoded version of the font. This is more or less a copy of the file OT1cmtt.fd.

The package accepts some options which may be useful:

override overrides the meaning of the \ttfamily command (and therefore also the \texttt command too), making it the same as the new \mttfamily command. This isn’t the default just in case the change breaks something in an unexpected way.

t1 informs the package that you’re using the T1 encoding, and therefore can borrow some accented characters from the DC version of Computer Modern Typewriter. This will probably be unnecessary, since the package attempts to work out what to do all by itself.

ot1 forces the package not to use the DC version of the Computer Modern Typewriter font for funny accents. Only use this option if the package thinks it should use the DC Typewriter font when it shouldn’t.

\mttfamily The command \mttfamily selects the properly-encoded Typewriter font. It’s a declaration which works just like the \ttfamily command, except that comamnds like \} and \_use the characters from the font rather than choosing odd-looking versions from the maths fonts. All of the accent commands still work properly. In fact, some accent commands which didn’t work before have been fixed. For example, saying ‘\texttt{P\'al Erd\H os}’ would produce something truly appalling like ‘Pál Erd} os’, which is obviously ghastly. The new encoding handles this properly, and produces ‘Pál Erdős’.^{1}

\textmtt Font changing commands are much more convenient than th declarations, so a command \textmtt is provided: it just typesets its argument in the re-encoded Typewriter font.

1This isn’t quite perfect. The accent, which isn’t actually present in the Typewriter font, is taken from the Computer Modern bold font, but it doesn’t look too bad. However, if you pass the option t1 to the cmtt package when you load it, the accent will be taken from the DC Typewriter font, and it will look totally wonderful.

\mtt Rather more excitingly, the \mtt command allows you to generate almostverbatim text very easily, without any of the restrictions of the \verb command. This command was inspired by something which David Carlisle said to me in an email correspondence regarding the overuse of verbatim commands.

\mtt redefines several ‘short’ commands to typeset the obvious characters. The complete list is shown below: there are some oddities, so watch out.

                  \\  \  \{  {  \}  }  \__\^^
                  \$  $  \%  %  \&  &  \#  #  \~  ~
                  \"  "  \'  '  \␣  ␣  \|  |

As well as redefining these commands, \mtt will endeavour to make single special characters display themselves in a verbatim-like way. This only works on ‘active’ characters (like ~), and \mtt makes no attempt to change the category codes of any characters.

Among other things, you’ll probably noticed that several accent-making commands have been redefined. You can still use these accents through the \a command, by saying \a', \a^and so on, as in the tabbing environment.

There are also some oddities in the table: | and " can be accessed easily without playing with silly commands. Well, that’s almost the case: these two characters are both often used as ‘short’ verbatim commands, so they are forced back to their normal meanings so you can type them.

Finally, a word on spacing. The \␣ command has been hijacked to produce a funny ‘visible space’ character. You can still produce multiple spaces by saying something like ‘␣{}␣{}… ␣’, which is a bit contrived, but that’s tough. Also, ~ has been stolen so that you can type ~ characters (e.g., in URLs), so the only way you can tpye a nonbreaking space is by using the \nobreakspace command, which is a bit of a mouthful. There’s an abbreviation, though: \nbsp now means exactly the same thing.

Was that not all supremely useful? Oh, just a note: this document doesn’t use a single verbatim command or environment (except in the listings, where it’s unavoidable) – it’s all done with \mtt.

2 Implementation

2.1 The package

1 ⟨∗sty⟩

I’ll start with some options handling.

 2 \newif\ifcmtt@override
 3 \newif\ifcmtt@dcfonts
 4 \def\@tempa{T1}\ifx\encodingdefault\@tempa
 5 \cmtt@dcfontstrue
 6 \fi
 7 \DeclareOption{override}{\cmtt@overridetrue}
 8 \DeclareOption{t1}{\cmtt@dcfontstrue}
 9 \DeclareOption{ot1}{\cmtt@dcfontsfalse}

10 \ProcessOptions

This bit is really trivial. I’ll just declare the font encoding. Oh, that was easy.

11 \DeclareFontEncoding{mTT}{}{}

Wait: there’s a problem. L^{A}TEX will now complain bitterly that it can’t find the font mTT/cmr/m/n, which is readonable, since I haven’t declared any such font. The following line should sort this out,

12 \DeclareFontSubstitution{mTT}{cmtt}{m}{n}

Now I’d better load all the text commands I’ll need when in this funny font variant.

13 \input{mTTenc.def}

\mttfamily \textmtt Finally, I’ll need to define a command which switches to this funny font, and a \text… command for it.

14 \DeclareRobustCommand{\mttfamily}{%
15 \fontencoding{mTT}\fontfamily{\ttdefault}\selectfont%

16 }

17 \DeclareTextFontCommand{\textmtt}{\mttfamily}

If an override was requested, make \ttfamily the same as \mttfamily.

18 \ifcmtt@override
19 \let\ttfamily\mttfamily
20 \fi

Well, that’s all that’s needed for the font definition. Here’s a command which will typeset its argument in the typewriter font, allowing easy access to all the funny characters, and printing them properly in the correct font (which \{ doesn’t do, for example).

\mtt@setchar This macro assigns the given meaning to the given control sequence. Also, if the character named in the control sequence is currently set active, it will set the active meaning of the character to the same value.

21 \def\mtt@setchar#1#2{%
22 \ifx#1#2\chardef#1`#1\else\let#1#2\fi%
23 \ifnum\catcode`#1=13%
24  \begingroup%
25  \lccode`\~=`#1%
26  \lowercase{\endgroup\let~#1}%
27 \fi%

28 }

\mtt@chars This macro lists the various control sequences which should be set up, so that they can be easily added to.

29 \def\mtt@chars{%
30 \do\#\#%
31 \do\%\%%
32 \do\&\&%
33 \do\^\^%
34 \do\~\~%
35 \do\'\textquotesingl%
36 \do\"\textquotedbl%
37 \do\|\textbar%

38 \do\$\textdollar%
39 \do\_\textunderscore%
40 \do\{\textbraceleft%
41 \do\}\textbraceright%
42 \do\\\textbackslash%
43 \do\ \textvisiblespace%

44 }

\mtt@do This just sets up all the special characters listed above. It’s a simple abbreviation, really.

45 \def\mtt@do{\let\do\mtt@setchar\mtt@chars}

\mtt And finally, the macro itself. Ta-da!

46 \DeclareRobustCommand\mtt[1]{\textmtt{\mtt@do#1}}

\@tabacckludge The otherwise almost totally perfect \@tabacckludge gets very upset when its argument is an active character. (If you’re wondering, this is the command which is responsible for the behaviour of the \a command.) Adding a \string makes everything work perfectly.

47 \def\@tabacckludge#1{%
48 \expandafter\@changed@cmd\csname\string#1\endcsname\relax%

49 }

50 \let\a\@tabacckludge

\nbsp Make an abbreviation for \nobreakspace.

51 \let\nbsp\nobreakspace

I think that’s all that I have to do for the package. If there’s any more to do, I’ll add it later.

52 ⟨/sty⟩

2.2 The font definition file

This is obviously copied almost verbatim from the file OT1cmtt.fd.

53 ⟨∗fd⟩

54 \DeclareFontFamily{mTT}{cmtt}{\hyphenchar\font\m@ne}
55 \DeclareFontShape{mTT}{cmtt}{m}{n}{
56 <5> <6> <7> <8> cmtt8
57 <9> cmtt9
58 <10> <10.95> cmtt10
59 <12> <14.4> <17.28> <20.74> <24.88> cmtt12
60 }{}
61 \DeclareFontShape{mTT}{cmtt}{m}{it}{
62 <5> <6> <7> <8> <9> <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>
63 cmitt10
64 }{}
65 \DeclareFontShape{mTT}{cmtt}{m}{sl}{
66 <5> <6> <7> <8> <9> <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>
67 cmsltt10
68 }{}
69 \DeclareFontShape{mTT}{cmtt}{m}{sc}{

70 <5> <6> <7> <8> <9> <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>
71 cmtcsc10
72 }{}
73 \DeclareFontShape{mTT}{cmtt}{m}{ui}  {<->sub * cmtt/m/it} {}
74 \DeclareFontShape{mTT}{cmtt}{bx}{n}  {<->sub * cmtt/m/n}  {}
75 \DeclareFontShape{mTT}{cmtt}{bx}{it} {<->sub * cmtt/m/it} {}
76 \DeclareFontShape{mTT}{cmtt}{bx}{ui} {<->sub * cmtt/m/it} {}

77 ⟨/fd⟩

2.3 The encoding definitions file

I’ve saved the trickiest bit until last. This file defines the mappings from text commands to glyphs in the font.

78 ⟨∗def⟩

First for some fun with accents. The cmtt font doesn’t contain all of the accents which the other Computer Modern fonts do, because those slots contain the standard ASCII characters which usually have to be ‘borrowed’ from the maths fonts.

Anyway, there’s a load which don’t need any special treatment. These are chosen from the OT1 encoding by default anyway, so I needn’t bother unless I’m really bothered about speed. I’m not, so I’ll save the memory.

Following the example of the TEXbook, I’ll use the bold roman font for accents, so that they don’t look really spindly. This is actually remarkably difficult to do, because the NFSS keeps getting in the way. I’ll look after the old font name in a macro (it’s handy that NFSS maintains this for me) and change to a known font, do the accent, change font back again, do the argument to the accent, and then close the group I did all of this in, so that no-one else notices what a naughty chap I am, really. This is startlingly evil.

79 \def\cmtt@accent#1#2{{%
80 \let\@old@font\font@name%
81 \ifcmtt@dcfonts%
82  \fontencoding{T1}\selectfont%
83 \else%
84  \usefont{OT1}{cmr}{bx}{n}%
85 \fi%
86 #1{\@old@font#2}%
87 }}

And now for the actual offending accents.

88 \DeclareTextCommand{\H}{mTT}{\cmtt@accent\H}
89 \DeclareTextCommand{\.}{mTT}{\cmtt@accent\.}

The ‘under’ accents are all OK, so I shan’t bother to define them either. Similarly, lots of the text symbol commands are fine as they are by default and I don’t need to try and define them again.

This, then, is the remaining commands which really need sorting out. (By the way, the only reason I’ve redefined \textellipsis is because otherwise it will mess up the nice monospacing.)

90 \DeclareTextSymbol{\textbackslash}{mTT}{92}
91 \DeclareTextSymbol{\textbar}{mTT}{124}
92 \DeclareTextSymbol{\textbraceleft}{mTT}{123}

93 \DeclareTextSymbol{\textbraceright}{mTT}{125}
94 \DeclareTextSymbol{\textless}{mTT}{60}
95 \DeclareTextSymbol{\textgreater}{mTT}{62}
96 \DeclareTextSymbol{\textunderscore}{mTT}{95}
97 \DeclareTextSymbol{\textvisiblespace}{mTT}{32}
98 \DeclareTextCommand{\textellipsis}{mTT}{...}
99 \DeclareTextSymbol{\textquotedbl}{mTT}{34}
100 \DeclareTextSymbol{\textquotesingl}{mTT}{13}

That’s all there is. Please return to your homes.

Appendix

A The GNU General Public Licence

The following is the text of the GNU General Public Licence, under the terms of which this software is distrubuted.

GNU GENERAL PUBLIC LICENSE Version 2, June 1991

Copyright (C) 1989, 1991 Free Software Foundation, Inc. 675 Mass Ave, Cambridge, MA 02139, USA Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

A.1 Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

A.2 Terms and conditions for copying, distribution and modification

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions:

(a) You must cause the modified files to carry prominent notices stating that you changed the files and the date of any change.

(b) You must cause any work that you distribute or publish, that in whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

(c) If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program.

In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following:

(a) Accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

(b) Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or, (c) Accompany it with the information you received as to the offer to dis-

tribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. Because the Program is licensed free of charge, there is no warranty for the Program, to the extent permitted by applicable law. except when otherwise stated in writing the copyright holders and/or other parties provide the program “as is” without warranty of any kind, either expressed or implied, including, but not limited to, the implied warranties of merchantability and fitness for a particular purpose. The entire risk as to the quality and performance of the Program is with you. Should the Program prove defective, you assume the cost of all necessary servicing, repair or correction.

12. In no event unless required by applicable law or agreed to in writing will any copyright holder, or any other party who may modify and/or redistribute the program as permitted above, be liable to you for damages, including any general, special, incidental or consequential damages arising out of the use or inability to use the program (including but not limited to loss of data or data being rendered inaccurate or losses sustained by you or third parties or a failure of the Program to operate with any other programs), even if such holder or other party has been advised of the possibility of such damages.

END OF TERMS AND CONDITIONS

A.3 Appendix: How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

<one line to give the program's name and a brief idea of what it does.>
Copyright (C) 19yy  <name of author>
This program is free software; you can redistribute it and/or modify
it under the terms of the GNU General Public License as published by
the Free Software Foundation; either version 2 of the License, or
(at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it

starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
This is free software, and you are welcome to redistribute it
under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items–whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in the program
`Gnomovision' (which makes passes at compilers) written by James Hacker.
<signature of Ty Coon>, 1 April 1989
Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\" . . . . . . . . . . . . . . . . . . . . . . . . . 36 \# . . . . . . . . . . . . . . . . . . . . . . . . . 30 \$ . . . . . . . . . . . . . . . . . . . . . . . . . 38 \% . . . . . . . . . . . . . . . . . . . . . . . . . 31 \& . . . . . . . . . . . . . . . . . . . . . . . . . 32 \' . . . . . . . . . . . . . . . . . . . . . . . . . 35 \. . . . . . . . . . . . . . . . . . . . . . . . . . 89 \@changed@cmd . . . . . . . . . . . . . . . . 48 \@old@font . . . . . . . . . . . . . . . . 80, 86 \@tabacckludge . . . . . . . . . . . . . . . \underline{47} \@tempa . . . . . . . . . . . . . . . . . . . . . . 4 \\ . . . . . . . . . . . . . . . . . . . . . . . . . 42 \{ . . . . . . . . . . . . . . . . . . . . . . . . . 40 \} . . . . . . . . . . . . . . . . . . . . . . . . . 41 \^. . . . . . . . . . . . . . . . . . . . . . . . . 33 \_. . . . . . . . . . . . . . . . . . . . . . . . . 39 \| . . . . . . . . . . . . . . . . . . . . . . . . . 37 \~ . . . . . . . . . . . . . . . . . . . . . . . 25, 34 \␣ . . . . . . . . . . . . . . . . . . . . . . . . . 43

A \a . . . . . . . . . . . . . . . . . . . . . . . . . 50

C \chardef . . . . . . . . . . . . . . . . . . . . 22 \cmtt@accent . . . . . . . . . . . . 79, 88, 89

                                       \cmtt@dcfontsfalse . . . . . . . . . . . . . 9
                                       \cmtt@dcfontstrue . . . . . . . . . . . 5, 8
                                       \cmtt@overridetrue . . . . . . . . . . . . . 7

D

                                       \DeclareFontEncoding . . . . . . . . . . 11
                                       \DeclareFontFamily . . . . . . . . . . . . 54
                                       \DeclareFontShape 55, 61, 65, 69, 73–76
                                       \DeclareFontSubstitution . . . . . . . 12

\DeclareOption . . . . . . . . . . . . . . . 7–9

                                       \DeclareRobustCommand . . . . . . . 14, 46
                                       \DeclareTextCommand . . . . . . 88, 89, 98

\DeclareTextFontCommand . . . . . . . 17
\DeclareTextSymbol . . . . 90–97, 99, 100

\do . . . . . . . . . . . . . . . . . . . 30–43, 45

E

\encodingdefault . . . . . . . . . . . . . . 4

F

\font . . . . . . . . . . . . . . . . . . . . . . . 54 \font@name . . . . . . . . . . . . . . . . . . 80 \fontencoding . . . . . . . . . . . . . . 15, 82 \fontfamily . . . . . . . . . . . . . . . . . . 15

H

\H . . . . . . . . . . . . . . . . . . . . . . . . . 88 \hyphenchar . . . . . . . . . . . . . . . . . . 54

I

\ifcmtt@dcfonts . . . . . . . . . . . . . 3, 81

\ifcmtt@override . . . . . . . . . . . . 2, 18

\input . . . . . . . . . . . . . . . . . . . . . . 13

L

\lccode . . . . . . . . . . . . . . . . . . . . . 25 \lowercase . . . . . . . . . . . . . . . . . . 26

M

\mtt . . . . . . . . . . . . . . . . . . . . . . 2, \underline{46} \mtt@chars . . . . . . . . . . . . . . . . \underline{29}, 45 \mtt@do . . . . . . . . . . . . . . . . . . . \underline{45}, 46 \mtt@setchar . . . . . . . . . . . . . . . \underline{21}, 45 \mttfamily . . . . . . . . . . . . . . 2, \underline{14}, 19

N

\nbsp . . . . . . . . . . . . . . . . . . . . . . . \underline{51} \newif . . . . . . . . . . . . . . . . . . . . . 2, 3 \nobreakspace . . . . . . . . . . . . . . . . 51

P

\ProcessOptions . . . . . . . . . . . . . . 10

S

\selectfont . . . . . . . . . . . . . . . . 15, 82

T

\textbackslash . . . . . . . . . . . . . 42, 90 \textbar . . . . . . . . . . . . . . . . . . 37, 91 \textbraceleft . . . . . . . . . . . . . 40, 92 \textbraceright . . . . . . . . . . . . 41, 93 \textdollar . . . . . . . . . . . . . . . . . . 38 \textellipsis . . . . . . . . . . . . . . . . 98 \textgreater . . . . . . . . . . . . . . . . . 95 \textless . . . . . . . . . . . . . . . . . . . 94 \textmtt . . . . . . . . . . . . . . . . 2, \underline{14}, 46 \textquotedbl . . . . . . . . . . . . . . 36, 99 \textquotesingl . . . . . . . . . . . 35, 100 \textunderscore . . . . . . . . . . . . 39, 96

\textvisiblespace . . . . . . . . . . 43, 97

\ttdefault . . . . . . . . . . . . . . . . . . 15 \ttfamily . . . . . . . . . . . . . . . . . . . 19

U

\usefont . . . . . . . . . . . . . . . . . . . . 84